I am a safety supervisor for an oil company. I need to know what is the force generated by an unsecured 300lb drum of oil traveling at 30mph if the truck stops suddenly. I am trying to instruct drivers about securing their loads and what can happen if they don't secure the load. For example: The 300lb drum would generate an impact force = to 1500lbs if the truck was traveling at 30 mph??? What about 50mph??? 2500 lbs of force????
Can you help me with these calculations?

Thanks,

Steve
1
Expert's answer
2019-03-13T12:41:48-0400

It's quite easy to calculate the force applying Newton's second law:


"F=ma"

and the formula for acceleration in terms of braking time and braking distance:


"a=\frac{2d}{t^2}."

Thus the impact force will be


"F=m\cdot \frac{2d}{t^2}."

On the other hand, you can use the same law but substitute the acceleration expressed in terms of speed and braking distance:


"F=m\cdot\frac{v^2_{final}-v^2_{initial}}{2d}."

Also, expressing the acceleration through velocities and the time required to slow the truck down from "v_{initial}" to "v_{final}":


"F=m\cdot\frac{v_{final}-v_{initial}}{t}."

Remember that to determine the force in Newtons, use m/s for speed, meters and seconds.